OpenStudy (anonymous):

I need to find any critical numbers of the function: h(x)=sin^2 x + cos x and 0

OpenStudy (anonymous):

find the first derivative and set it =0 :)

OpenStudy (anonymous):

lol i got that. it was the finding the derivative part that sucks.

OpenStudy (anonymous):

h(x)=sin^2 x + cos x and 0<2pii Dh(x)=D(sin^2 x + cos x) =Dsin^2 x + Dcos x let u=sinx so that Du^2=2uDu therefore =2sin x cosx -sinx=sin 2x -sinx